### v2

# Learning Optimal Stable Matches in Decentralized Markets with Unknown Preferences

Abstract
Matching algorithms have demonstrated great success in several practical applications, but they often require centralized coordination and plentiful information. In many modern online marketplaces, agents must independently seek out and match with another using little to no information. For these kinds of settings, can we design decentralized, limited-information matching algorithms that preserve the desirable properties of standard centralized techniques? In this work, we constructively answer this question in the affirmative. We model a two-sided matching market as a game consisting of two disjoint sets of agents, referred to as proposers and acceptors, each of whom seeks to match with their most preferable partner on the opposite side of the market. However, each proposer has no knowledge of their own preferences, so they must learn their preferences while forming matches in the market. We present a simple online learning rule that guarantees a strong notion of probabilistic convergence to the welfare-maximizing equilibrium of the game, referred to as the proposer-optimal stable match. To the best of our knowledge, this represents the first completely decoupled, communication-free algorithm that guarantees probabilistic convergence to an optimal stable match, irrespective of the structure of the matching market.
I Introduction
How should partnerships be formed? Although ostensibly simple, the task of matching members of two distinct groups is inherently combinatorially challenging. The two-sided matching problem has thus motivated decades of research by computer scientists, economists, and engineers alike that has fundamentally transformed the landscape of our social and engineered systems [1, 2, 3, 4].
Consider the residency matching process, a prototypical example of a two-sided matching problem in which residents and hospitals must be partnered with one another. In practice, residents typically rank order the hospitals, as do hospitals rank order the residents. A central matchmaker would ideally use these orderings to identify a desirable match between residents and hospitals; in particular, one would hope to find a stable match, characterized by the property that no resident and hospital would mutually prefer one another to their partners assigned by the matchmaker.
Celebrated work by Gale and Shapley [5] demonstrates the existence of such a stable match in every two-sided market. Given a list of the preference orderings of every agent, the Gale-Shapley algorithm (GS) identifies the best possible stable match for all agents on one side of the market (Figure 1, top-left). GS and its variants have been implemented in residency admissions and organ donation programs, guaranteeing strong notions of stability and optimality to the arrangements made by these institutions [6, 7].
However, GS is not applicable in all two-sided matching settings. In many scenarios, there is no central clearinghouse that facilitates the matching process, so an algorithm like GS cannot be used. In these kinds of markets (e.g., most labor markets), there is typically ‘active’ set of agents, the proposers (firms), who repeatedly make proposals to a ‘passive’ set of agents, the acceptors (workers), who accept or reject these proposals (Figure 1, top-right). Motivated by these kinds of decentralized environments, scholars have extensively studied the stability and quality of the matches that may result when agents match independently [8, 9, 10].
Figure 1: Two-sided matching market models with full (top) and partial information (bottom) in centralized (left) and decentralized environments (right). In a full information model, proposers know their own preferences completely, but in a partial information model, they do not. In a centralized environment, agents provide their preferences (or a subset thereof) to a central algorithm that assigns a match, but in a decentralized environment, proposers directly propose to acceptors to form a match.
GS and similar algorithms are well-suited for situations where users report comprehensive preference orderings, but there are many applications where such requirements are prohibitive (Figure 1, bottom-left). Generating rankings may become difficult because of issues related to size, a lack of information, or both. To address these issues, large practical markets like the National Resident Matching Program [6, 7, 11] use specialized algorithms designed to handle shortened or nonexistent preference orderings [12, 13].
This work focuses on markets that face a lack of both centrality and information. A particular motivation is the modern online marketplace, where massive numbers of consumers and producers seek to exchange resources without information on one another or guidance from a central entity (Figure 1, bottom-right). Agents repeatedly interact with one another and receive only limited information on the quality of their matched partners through the form of some representative reward. In these kinds of settings, there is a growing need for decentralized learning rules, policies that allow agents to learn from their past rewards so that their future behavior leads towards collectively desirable stable matches.
Several recent papers have made significant progress towards this goal for the setting where agents on the proposing side of the market do not know their own preferences [14, 15]. Many draw from the multi-armed bandits literature to construct algorithms that provide bounds on the regret experienced by each proposer [16, 17, 18, 19, 20], while others use more general probabilistic techniques that yield similar results [21, 22]. However, these learning rules typically require some form of online communication or coordination between proposers or a central entity, and they also impose restrictive assumptions on the structure of the agents’ preferences. Furthermore, the guarantees on the stability or quality of the resulting matches when using these rules are limited.
In this paper, we present a novel algorithm that guarantees stability in decentralized partial-information settings. In particular, we design a learning rule for the proposers to independently follow that provably guarantees a strong notion of probabilistic convergence to their optimal stable match even when they do not know their own preferences. The learning rule thus recovers the classic result of Gale and Shapley for the case of one-sided unknown preferences.
Our learning rule draws from the literature on learning in games, a well-established subfield of game theory concerned with the design of online learning algorithms [23, 24]. For several types of games, scholars have designed decentralized learning rules that guarantee convergence to desirable configurations such as correlated, Nash, and even welfare-maximizing equilibria [25, 26, 27, 28, 29]. Inspired by these results, we first formulate the decentralized two-sided matching market as a game in Section II. Then, in Section III, we assert the existence of a novel learning rule that guarantees convergence to the proposer-optimal stable match. We describe this rule in Section IV and prove its convergence properties in Section VI. Simulations and additional proof details are available in the full version of the text [30].
II Model
II-A Market Model
The two-sided market consists of two groups: a set of proposers, denoted by $\mathbf{P}\triangleq\{P_{1},P_{2},\dots,P_{n}\}$, and a disjoint iset of acceptors, denoted by $\mathbf{A}\triangleq\{A_{1},A_{2},\dots,A_{m}\}$. Each agent has a complete, transitive, and strict preference ordering, also referred to as preferences, over the agents in the other group. We represent the preference ordering of proposer $P_{i}$ as a function [^1] We choose this interval for simplicity, but any bounded interval (open or closed) is acceptable. $O_{P_{i}}:\mathbf{A}\cup\{\emptyset\}\to[0,1]$, where we say that $P_{i}$ prefers $A_{j_{1}}$ to $A_{j_{2}}$ if $O_{P_{i}}(A_{j_{1}})>O_{P_{i}}(A_{j_{2}})$, which we shorthand as $A_{j_{1}}\succ_{i}A_{j_{2}}$; for completeness, we assume $O_{P_{i}}(A_{j})>O_{P_{i}}(\emptyset)=0$ for all $j\in\{1,\dots m\}$. We define similarly the preference ordering $O_{A_{j}}$ of acceptor $A_{j}$, and as a slight abuse of notation, we use the same shorthand $\succ_{j}$ to describe the preferences of acceptor $A_{j}$. The collection of all preference orderings is given by $\mathbf{O}\triangleq\{O_{P_{1}},\dots,O_{P_{n}},O_{A_{1}},\dots,O_{A_{m}}\}$, so that a particular instance of the matching market game is fully parametrized by the tuple $\mathbf{M}=(\mathbf{P},\mathbf{A},\mathbf{O})$. A simple market is depicted in Figure 2 with $n=m=3$ proposers and acceptors.
Figure 2: A market with three proposers (left) and three acceptors (right). Each agents’ preference ordering is shown as a list beneath their icon. The arrows indicate to whom each proposer proposes; this configuration corresponds to the proposer-optimal stable match in which every proposer is matched with their favorite acceptor.
II-B Game Model
We study interactions between agents in the two-sided matching market using a game theoretic framework. A market $\mathbf{M}$ defines a game played by the set of proposers, where each proposer $P_{i}\in\mathbf{P}$ selects an action $a_{i}\in\mathbf{A}\cup\{\emptyset\}$. If $a_{i}=A_{j}$, we say that $P_{i}$ proposes to $A_{j}$, and if $a_{i}=\emptyset$, we say that $P_{i}$ chooses to remain single. The full action profile is denoted by $a\triangleq\{a_{1},\dots,a_{n}\}$. After the proposers propose, each acceptor $A_{j}$ accepts their most preferred proposer in $\{P_{i}\in\mathbf{P}:a_{i}=A_{j}\}$ (where ‘most preferred’ is specified by $O_{A_{j}}$), and rejects all other proposers that have proposed to them. Since we assume that acceptors behave determinstically by accepting their most preferred proposal from those they receive, we do not regard them as players in the game.
We denote the outcome of these acceptances and rejections as a match $\mu(a)\triangleq\{\mu_{P_{1}}(a),\dots,\mu_{P_{n}}(a),\mu_{A_{1}}(a),\dots,\mu_{A_{m}}(a)\}$, where $\mu_{P_{i}}(a)\in\mathbf{A}\cup\{\emptyset\}$ is the match (or lack thereof) of proposer $P_{i}$ given the action profile $a$, and $\mu_{A_{j}}(a)\in\mathbf{P}\cup\{\emptyset\}$ denotes the outcome for $A_{j}$. If $\mu_{P_{i}}(a)=A_{j}$ then $\mu_{A_{j}}(a)=P_{i}$, and we say that the pair $(P_{i},A_{j})$ are matched. Otherwise, if $\mu_{P_{i}}(a)=\emptyset$ or $\mu_{A_{j}}(a)=\emptyset$, then we say that proposer $P_{i}$ or acceptor $A_{j}$ is unmatched. For ease of notation, we simply write $\mu$, $\mu_{P_{i}}$ or $\mu_{A_{j}}$ when the dependence on $a$ is clear.
Using this notation, we can define the utility for each proposer $P_{i}$ as
$u_{i}(a)=O_{P_{i}}(\mu_{P_{i}}(a)),$
which is the payoff they receive for matching with their partner. With this definition, the assumption $O_{P_{i}}(\emptyset)=0$ ensures that every proposer strictly prefers being matched to being unmatched.
Let us now discuss the equilibria of the game. Any match $\mu$ in which there exists a proposer-acceptor pair $(P_{i},A_{j})$ who prefer one another to their partners in $\mu$ (i.e., $O_{P_{i}}(A_{j})>O_{P_{i}}(\mu_{P_{i}})$ and $O_{A_{j}}(P_{i})>O_{A_{j}}(\mu_{A_{j}})$) is said to be unstable, and any match that is not unstable is said to be stable. A stable match exists in every two-sided market [5]. Moreover, there exists a unique proposer-optimal stable match (POSM) $\mu^{*}$, which is characterized by the property that, for any alternate stable match $\mu^{\prime}$ in the market,
$\mu^{*}_{P_{i}}\succeq_{i}\mu^{\prime}_{P_{i}}\qquad\forall i\in\{1,\dots,n\},$
meaning that every proposer weakly prefers their partner in the POSM to their partner in any other stable match. For example, the proposer optimal stable match in the market shown in Figure 2 can be described by the list of proposer-acceptor pairs $(P_{1},A_{1}),(P_{2},A_{2}),(P_{3},A_{3})$.
From this, one can conclude that every pure strategy Nash equilibrium of the game corresponds to a stable matching. More specifically, every Nash equilibrium action profile produces exactly one stable match; however, a stable match may be represented by multiple Nash equilibrium action profiles if $n>m$, since the proposers who are unmatched in a stable match are indifferent between their actions. Furthermore, if one defines the welfare of an action profile as the sum of the players’ utilities corresponding to that action profile, then it is straightforward to show that any action profile corresponding to the proposer-optimal stable match is a welfare-maximizing Nash equilibrium of the game.
III Learning Model and Main Result
We model the repeated interactions between proposers and acceptors as a series of repeated one-shot matching games that are played in timesteps $t=\{1,2,\dots\}$. At timestep $t$, every proposer $P_{i}$ selects their action $a_{i}^{t}$, which generates a match $\mu^{t}$ and a utility $u_{i}^{t}(a^{t})$ as described in the previous section. We assume that proposer $P_{i}$ selects $a_{i}^{t}$ according to a probability distribution $p_{i}^{t}$, which we refer to as their strategy. By assumption, their strategy at timestep $t$ can only incorporate their actions and utilities in previous timesteps $11,2,\dots,\;t-$ 1. We thus define a learning rule as a mechanism that agents use to generate their strategy at each timestep, i.e.,
$p_{i}^{t}=\Gamma(a_{i}^{1},u_{i}^{1},\dots,a_{i}^{t-1},u_{i}^{t-1};\mathbf{A}\cup\{\emptyset\}).$ (1)
Notice that when using $\Gamma$, a proposer must know only their own action and utility history, as well as their own action space, but nothing else. In particular, they need not have any knowledge of their own true preference ordering, nor do they require information regarding the existence of other proposers or their actions. Such learning rules are described as completely uncoupled.
Having defined the relevant concepts, let us now formally restate the central question studied in this paper:
Is there a learning rule that provably guarantees convergence to the proposer-optimal stable match in decentralized markets where proposers do not know their own preferences?
We assert the existence of a novel learning rule that answers this question in the affirmative in Theorem 1.
Theorem 1.
Consider the matching game defined by the market $\mathbf{M}$. For every probability $10\leq p<$ 1, there exists a learning rule $\Gamma$ of the form defined in (1) such that, if every proposer selects their action according to $\Gamma$, then for all sufficiently large timesteps $t$, the action profile $a^{t}$ corresponds to the proposer-optimal stable match $\mu^{*}$ with probability at least $p$.
We take this opportunity to highlight the generality of this result. That is, Theorem 1 holds true without any assumptions on the number of proposers and acceptors or the structure of the preference orderings (i.e., the sequential preference condition, $\alpha$-reducibility), which are necessary for many previous results [16, 17, 18]. Furthermore, applying the rule is infrastructurally unrestrictive, in the sense that agents must only be informed of $\Gamma$ prior to the game being played; otherwise, there is no need for any platform that facilitates the exchange of information between agents or a central entity for the remainder of the process. This is a necessity in multiple existing methods [19, 20] that is not always available in practice.
The specific structure of the learning rule that achieves the desired result in Theorem 1 is discussed is detail in Section IV. However, we want to emphasize that we do not view the specifics of this learning rule as a key contribution of this paper. Rather, the main contribution resides in the fact that one can achieve the POSM in fully distributed settings with almost no information.
Lastly, we note that the proof of our result requires building upon a number of results on better- and best-reply dynamics in matching games [31, 32], which is quite involved. We argue that both the result and its proof present interesting contributions to matching theory.
IV Description of Learning Rule
In this section, we define the learning rule that each agent uses to select their action, which follows a similar structure [^2] In [28], the authors present a learning rule that guarantees convergence to welfare-maximizing equilibria. However, these dynamics only apply to games that are interdependent (see [28]), which the matching game is not. to the trial and error learning dynamics considered in [24, 27, 28]. In the learning rule, each proposer maintains a local state variable that impacts their chosen strategy at each timestep. After implementing the given strategy, they receive a utility that in turn influences their state, and the feedback loop between these two processes continues. The specific rules that govern how proposers select their actions and update their states are spelled out in detail below.
In each timestep, proposer $P_{i}$ generates their strategy $p_{i}^{t}$ according to their state $x_{i}\triangleq(m_{i},\underline{a}_{i},\underline{u}_{i})$, where
  • $m_{i}\in\{C,D,W\}$ is the mood of proposer $P_{i}$, which can take one of three values: $C$ (content), $W$ (watchful), or $D$ (discontent);
  • $\underline{a}_{i}\in\mathbf{A}\cup\{P_{i}\}$ is the baseline action of proposer $P_{i}$;
  • $\underline{u}_{i}\in\mathbb{R}_{\geq 0}$ is the baseline utility of proposer $P_{i}$.
We write the collection of proposer states as $x=(m,\underline{a},\underline{u})$. The time-homogeneous processes by which proposers select their actions and update their states are referred to as the action selection rule and the state update rule, both of which are governed by a common experimentation rate $\epsilon>0$; we refer to the pair as the learning rule. At each timestep, every proposer independently selects and plays their action according to the action selection rule described in Algorithm 1.
Algorithm 1 Action selection rule
$P_{i}$, $x_{i}$, $\epsilon$
if $m_{i}=C$ then
     play baseline action $a_{i}=\underline{a}_{i}$ with probability (w.p.) $1-\epsilon-\epsilon^{2}$; or
     play new action $a_{i}\in\mathbf{A}$ uniformly at random w.p. $\epsilon$; or
     play new action $a_{i}=P_{i}$ w.p. $\epsilon^{2}$
else if $m_{i}=D$ then
     play baseline action $a_{i}=\underline{a}_{i}=P_{i}$ w.p. $1-\epsilon-\epsilon^{1.5}$; or
     play new action $a_{i}\in\mathbf{A}$ uniformly at random w.p. $\epsilon^{1.5}$
else if $m_{i}=W$ then
     play baseline action $a_{i}=\underline{a}_{i}$ w.p. $1$
After selecting and playing their action $a_{i}$ and receiving their utility $u_{i}$, each proposer $P_{i}$ independently updates their state $x_{i}$. This update behavior is governed by two functions $F:[0,1]\to[0,0.5)$ and $G:[0,1]\to[0,0.5)$, where we assume that both functions are strictly monotone decreasing [^3] The domains and codomains are chosen for simplicity of analysis, but other intervals can be used with appropriate modifications to the model.. The state update rule is described in Algorithm 2, where $\Delta u_{i}\triangleq u_{i}-\underline{u}_{i}$.
Algorithm 2 State update rule
$P_{i}$, $x_{i}$, $u_{i}$, $\epsilon$, $F$, $G$
if $m_{i}=C$ and $P_{i}$ did not experiment then
     $x_{i}=\begin{cases}(C,\underline{a}_{i},u_{i})&u_{i}\geq\underline{u}_{i}\\ (W,\underline{a}_{i},\underline{u}_{i})&u_{i}<\underline{u}_{i}\end{cases}$
else if $m_{i}=C$ and $P_{i}$ did experiment then
     if $u_{i}\leq\underline{u}_{i}$ then
         $x_{i}=(C,\underline{a}_{i},u_{i})$
     else
         $x_{i}=\begin{cases}(C,\underline{a}_{i},\underline{u}_{i})&\text{w.p. }1-\epsilon^{G(\Delta u_{i})}\\ (C,a_{i},u_{i})&\text{w.p. }\epsilon^{G(\Delta u_{i})}\end{cases}$
else if $m_{i}=D$ and $P_{i}$ did experiment then
     $x_{i}=\begin{cases}(D,\underline{a}_{i},\underline{u}_{i})&\text{w.p. }1-\epsilon^{F(u_{i})}\\ (C,a_{i},u_{i})&\text{w.p. }\epsilon^{F(u_{i})}\end{cases}$
else if $m_{i}=W$ then
     $x_{i}=\begin{cases}(C,\underline{a}_{i},\underline{u}_{i})&u_{i}\geq\underline{u}_{i}\\ (D,P_{i},0)&u_{i}<\underline{u}_{i}\end{cases}$
Although lengthy, the rule prescribes qualitatively straightforward behavior. Proposers asynchronously alternate between a content mood, where they successfully match in the market, and a discontent mood, where they are inactive and remain single. Occasionally, the learning rule spurs them to experiment with new actions, which probabilistically induces changes in their own state as well as others’ depending on changes in utility.
V Conclusion
In this work, we study a two-sided matching market in which the agents on the proposing side of the market do not know their preferences over the agents on the accepting side. We provide a fully decentralized learning rule that, if every agent follows, guarantees a strong notion of probabilistic convergence to the proposer-optimal stable match. In doing so, we recover the result of Gale and Shapley for decentralized markets with one-sided unknown preferences.
This work closes a gap in the literature on matching markets with unknown preferences. In particular, to the best of the authors’ knowledge, the learning rule presented in this paper is the first to provably guarantee any notion of convergence to a stable match, much less the optimal stable match, and does so without any restrictions on the agents’ preference orderings. We view our learning rule as a theoretically grounded starting point for developing more advanced algorithms that increase the rate of convergence while maintaining provable guarantees.
VI Proof of Theorem
The proof of Theorem 1 relies on several existing results from the literature on learning in games and two-sided matching markets, as well as some novel results on best-response dynamics in matching games available in the full version of the text [30]. We combine these results to complete our proof below.
Observe that the learning rule induces a regularly perturbed Markov process over the finite state space
$\mathbf{X}\triangleq\prod_{i=1}^{n}\{C,D,W\}\times(\mathbf{A}\cup\{P_{i}\}\times(\text{ran}(O_{P_{i}})\cup\{0\}),$
where $\text{ran}(f)$ denotes the range of the function $f$. We must show that the stochastically stable states of this process correspond to the POSM [23, Theorem 4] We write the probability transition matrix over the state space $\mathbf{X}$ as $\mathbf{T}^{\epsilon}$ for each $\epsilon>0$. We must first identify the recurrence classes of the unperturbed process $\mathbf{T}^{0}$. Let $\mathbf{Z}\subseteq\mathbf{X}$ denote the subset of states in which every agents’ benchmark utilities are aligned with their benchmark actions, i.e., no agent is watchful. Let $\mathbf{C}$ denote the subset of states in which at least one matched (content) proposer is not playing a best response, and let $\mathbf{D}$ denote the subset of states in which every matched proposer is playing a best response, but there is at least one unmatched (discontent) proposer. Let $\mathbf{E}\subset\mathbf{Z}$ denote the subset of states whose benchmark action profiles correspond to stable matches, and let $e^{*}\in\mathbf{E}$ denote the state whose benchmark action profile correspond to the POSM. Observe that $\mathbf{C}$, $\mathbf{D}$, and $\mathbf{E}$ form a partition of $\mathbf{Z}$.
The recurrence classes of the unperturbed process are all singletons $\{x\}$ such that $x\in\mathbf{Z}$. To see this, first consider a state in which every agent is not watchful. Observe that in the unperturbed process, no one experiments, so every content and discontent proposer will repeat their baseline action and receive their same baseline utility. Thus, the process remains in state $x$ with probability one. Next, consider a state in which one more agents are watchful. Once again, every proposer will repeat their baseline action, and after receiving their utility, every watchful proposer will become content or discontent with probability one. Hence their benchmarks will align, meaning that the process arrives at a state in $\mathbf{Z}$ in exactly one transition.
Next, we introduce some important definitions that will be essential for the remainder of our proof. Define the function $r^{*}:\mathbf{Z}\to\mathbb{R}$ as follows:
$r^{*}(z)\triangleq\min\big{\{}r(z,z^{\prime})\big{|}\;z^{\prime}\in\mathbf{Z}\setminus\{z\}\big{\}}.$
Let $G$ be the graph constructed from $\mathbf{Z}$ as described in Section VII-A in [30]. We say that the edge from $z$ to $z^{\prime}$ is easy if $r(z,z^{\prime})=r^{*}(z)$. Similarly, we say that a path or tree is easy if each of its constituent edges are easy. The remainder of the proof involves showing that easy paths leads to the POSM, implying that the only easy tree is rooted at the POSM. We begin by identifying the nature of easy transitions for various states.
First, consider a state $z\in\mathbf{C}$, and suppose that the baseline action profile is such that at least one content proposer is not playing a best response. We claim that any easy edge from $z$ to some other state $z^{\prime}$ must involve a content proposer updating to a best response action. To see this, observe that the only possible easy transitions out of $z$ require experimentation. In particular, proposers can induce a transition out of $z$ by experimenting twice in succession without updating their baseline actions, causing at least one proposer to first become watchful and then discontent. Every other possible easy transition out of $z$ requires a proposer to experiment and update their baseline action: the resistance of a transition whereby a content proposer chooses to remain single is $2$, the resistance of a transition whereby a discontent proposer chooses to experiment and update their action is $1.5+F(\Delta u)$, and the resistance of a transition whereby a content proposer chooses to experiment and update their action is $1+G(\Delta u)$, where $\Delta u$ is the change in utility that the updating proposer experiences. Since $0.50\leq F(\cdot),G(\cdot)<$ 0.5, it is clear that a transition of minimum resistance must involve a content proposer experimenting and updating their action; moreover, since $G$ is strictly monotone decreasing, this update must be a best response. Given that no one experiments in the following timestep, which occurs with probability $O(1)$, the process transitions to a state $z^{\prime}\in\mathbf{Z}$ in two timesteps, and the total resistance along this path is $1+G(\Delta u)$.
Next, consider a state $z\in\mathbf{D}$, and suppose that the baseline action profile is such that every content proposer is playing a best response, and there is at least one discontent proposer. We claim that any easy edge from $z$ to some other state $z^{\prime}$ must involve a discontent proposer updating to a best response action. To see this, first notice that since content proposers are already playing best responses, they cannot unilaterally deviate to improve their utilities. This implies that any transition out of $z$ in which a content proposer experiments must involve at least two simultaneous or consecutive experimentations, which would have resistance at least $2$. Hence, a transition of minimum resistance must involve a discontent proposer experimenting and updating their action; moreover, since $F$ is strictly monotone decreasing, this update must be a best response. Given that no one experiments in the following timestep, which occurs with probability $O(1)$, the process transitions to a state $z^{\prime}\in\mathbf{Z}$ in two timesteps, and the total resistance along this path is $1+F(\Delta u)$, which is strictly less than $2$.
Lastly, consider a state $e\in\mathbf{E}$. We claim that $r^{*}(e)=2$, and that any easy transition out of $e$ can only occur in one of two ways. First, observe that since $e$ is a stable match, any unmatched proposer cannot unilaterally deviate to affect the outcome of any agent, including themselves. Hence, an easy transition will not involve a discontent proposer. However, by the same argument as before, a content proposer can induce a transition out of $z$ with resistance $2$. In particular, two consecutive experimentations can cause a transition to a state $z^{\prime}$ in which the baseline action profile is a near-stable match. However, this form of transition may not be possible depending on the preference structure $\mathbf{O}$. The other form of easy transition out of $z$ with resistance $2$ involves a content proposer choosing to remain single, which causes a one-timestep transition to a state $z^{\prime}$ in which the baseline action profile is a near-stable match. This form of transition is always possible.
The preceding arguments imply that every easy path starting from a state $z\notin\mathbf{E}$ consists of a sequence of baseline action profiles which are generated by the best response dynamics. In particular, an easy transition out of a state in $\mathbf{C}$ involves a matched proposer playing a best response, which occurs in Phase 1 of the best response dynamics, and an easy transition out of a state in $\mathbf{D}$ involves an unmatched proposer playing a best response, which occurs in Phase 2 of the best response dynamics.
Having established the nature of these easy paths, we can now provide a simple construction of an easy tree rooted at $e^{*}$. Initialize $T(e^{*})$ as an empty set of edges, and add to it as follows. Consider a vertex $z_{1}\in\mathbf{Z}\setminus\mathbf{E}$. If $z_{1}$ is not part of an edge already in $T(e^{*})$, then add the easy edge from $z_{1}$ to $z_{2}$. Continue adding edges in this fashion, traversing nodes $z_{3},\dots,z_{k}$ until either $z_{k}$ already has an outgoing edge in $T(e^{*})$, or until reaching an equilibrium $z_{k}=e_{1}\in\mathbf{E}$. Repeat this process until every node in $\mathbf{Z}\setminus\mathbf{E}$ has exactly one outgoing edge. Then, for every vertex $e\in\mathbf{E}\setminus\{e^{*}\}$, add an easy edge from $e$ to a state $z\in\mathbf{Z}$ whose baseline action constitutes a near-stable match; in particular, choose a state such that the proposer made single belongs to a rotation with respect to the stable match represented by $e_{1}$. However, from [30, Lemma 3], the easy path from $e$ will lead to an equilibrium that every proposer weakly prefers to a next-stable match with respect to the stable match given by $e$, because a proposer in a rotation is made single. It is straightforward to show that the properties of the best response dynamics ensure that this construction does not introduce any cycles. Thus, there exists a unique easy path from every vertex $z\in\mathbf{Z}\setminus\{e^{*}\}$ to $e^{*}$ in $T(e^{*})$, so $T(e^{*})$ is indeed an easy tree rooted at $e^{*}$.
All that remains to be shown is that a minimum resistance tree rooted at any other node $z\neq e^{*}$ has resistance strictly greater than $\rho(e^{*})$, the stochastic potential of $e^{*}$. First, consider any $z\in\mathbf{C}\cup\mathbf{D}$, and let $T(z)$ be a minimum resistance tree rooted at $z$. Every vertex $z^{\prime}\in\mathbf{Z}\setminus\{z,e^{*}\}$ contributes resistance at least $r^{*}(z^{\prime})$, the vertex $e^{*}$ contributes resistance at least $2$, and the resistance corresponding to $z$ is removed, so the stochastic potential of $T(z)$ is at least $\rho(z)\geq\rho(e^{*})+2-r^{*}(z)$. Because $z$ is not an equilibrium, $r^{*}(z)$ is strictly less than $2$, so $\rho(z)>\rho(e^{*})$. Next, consider a state $e\in\mathbf{E}$. Since $r^{*}(e)=r^{*}(e^{*})=2$, we have that $\rho(e)\geq\rho(e^{*})$. Because $T(e)$ is a tree rooted at $e$, there exists a unique path from $e$ to $e^{*}$. However, [30, Corollary 1] implies that every best response sequence starting from a near-stable match with respect to $\mu^{*}$ must lead back to $\mu^{*}$. Thus, the path from $e$ to $e^{*}$ in $T(e)$ must include at least one transition, say, from $z$ to $z^{\prime}$, which is not a best response update (if $z\in\mathbf{C}\cup\mathbf{D}$) or does not lead to a near-stable match (if $z\in\mathbf{E}$). This transition is not easy, so it has resistance strictly greater than $r^{*}(z)$; thus, $\rho(e)>\rho(e^{*})$. This establishes that the unique recurrence class of minimum stochastic potential is $e^{*}$, and hence, the POSM is the only stochastically stable state.